OpenStudy (anonymous):

Can I get some help Simplifying complex fraction (Algebra 2)?

OpenStudy (anonymous):

Simplify \[\frac{ 18x-6 }{ 9x^5 }\div \frac{ 15x+5 }{ 21x^2 }\]

Nnesha (nnesha):

you can change division to multiplication \[\large\rm \frac{a}{b} \div \frac{c}{d} = \frac{a}{b} *\frac{d}{c}\]multiply the first fraction by the reciprocal of the 2nd fraction

OpenStudy (anonymous):

Oh ok I'm working on the now give me second

OpenStudy (anonymous):

\[\frac{ 18x-6 }{ 9x5 }\times \frac{ 21x^2 }{ 15x+5 }\] does this look right?

Nnesha (nnesha):

yes correct now factor the binomial

OpenStudy (anonymous):

both of them?

Nnesha (nnesha):

yes.

OpenStudy (anonymous):

\[\frac{ 6(3x-1) }{ 9x^5 }\times \frac{ 21x^2 }{ 5(3x-1) }\] is this right?

OpenStudy (anonymous):

i think the next step crossing out the 3x-1 since they both have that in common?

Nnesha (nnesha):

that's correct!

OpenStudy (anonymous):

So now i have \[\frac{ 6 }{ 9x^5 }\times \frac{ 21x^2 }{ 5 }\]

Nnesha (nnesha):

good what else you can do to simplify ?

OpenStudy (anonymous):

get rid of the x's?

Nnesha (nnesha):

how would you do that ?

OpenStudy (mathmale):

Look for ways to reduce this expression further. Note that 6, 21 and 9 share a common factor. Also note that \[\frac{ x^2 }{ x^5 }\] can be reduced.

OpenStudy (anonymous):

\[\frac{ 2 }{ 3x^3 }\times \frac{ 7 }{ 5 }\]

OpenStudy (anonymous):

is this right?

OpenStudy (mathmale):

Could you simplify that result further, by multiplying as indicated?

Nnesha (nnesha):

okay so try to write each step thats way you will stay away from any kind of mistake \[\frac{6 }{ \color{ReD}{3} { \cancel{9}}x^5 }\times \frac{\color{ReD}{ 7} {\cancel{21}}x^2 }{ 5 }\] now how would you reduce the first fraction ?

OpenStudy (anonymous):

\[\frac{ 2 }{ 1x^5 }?\]

OpenStudy (anonymous):

is this correct?

Nnesha (nnesha):

that's right 6 divided by 3 would be 2/1 \[\frac{\color{blue}{2}{\cancel{6}} }{ \color{blue}{1}{\cancel{3}} { \cancel{9}}x^5 }\times \frac{\color{ReD}{ 7} {\cancel{21}}x^2 }{ 5 }\] \[\rm \frac{ 2 }{ 1x^3 } *\frac{7}{5}\]

OpenStudy (anonymous):

what do i do from this step?

Nnesha (nnesha):

is there anything else you can cancel out ??

OpenStudy (anonymous):

I don't think so?

Nnesha (nnesha):

alright thats it now you can just multiply

OpenStudy (anonymous):

would i get \[\frac{ 10 }{ 7x^3 }\]

OpenStudy (anonymous):

WAIT WAIT

OpenStudy (anonymous):

I SEE MY MISTAKE

Nnesha (nnesha):

good lol

OpenStudy (anonymous):

\[\frac{ 14 }{ 5x^3 }\] should be the answer

Nnesha (nnesha):

ye that's correct !! ^^ \[\rm \frac{ a }{ b } *\frac{c}{d} = \frac{a*c}{b*d}\]

Nnesha (nnesha):

good job! :=))

OpenStudy (anonymous):

HEYYOOO